# Resolving space-time structures of quantum impurities with a numerically exact few-body algorithm

### v1

@@ Abstract
Taking advantage of time-dependent rotations of electronic orbitals, we introduce a numerically exact real-time evolution scheme for quantum impurities in a macroscopic bath, revealing that almost all rotated orbitals can be made inactive (in a trivial product state). Following a quench, we show that both the number of active orbitals and their associated matrix product state bond dimension saturate to small values, leading to an algorithm dramatically more accurate and faster than the state of the art. We are thus able to follow precisely the dynamics for thousands of fermions, up to the long-time stationary regime, and to study subtle aspects of quantum relaxation in the temporal and spatial domain, such as the emergence of entanglement structures in the Kondo screening cloud.
@@
The many-body dynamics of generic quantum systems currently challenges numerical simulations, as exemplified by the unbounded entanglement growth in time of matrix product states (MPS). This situation also plagues quantum impurity problems, “simpler” systems still comprised of a macroscopic number of particles (bosons or fermions), where the interaction is restricted to a subset of degrees of freedom [1]. The interest in impurity models, originating in Kondo physics, has not waned over the years, since they are the cornerstone of the Dynamical Mean-Field Theory (DMFT) for correlated electrons [2]. Open quantum systems, in which qubits exchange energy and information with an environment, constitute also a very important class of impurity models [3]. However, in most cases, simulating the real-time non-perturbative dynamics of impurity models comes with tremendous numerical cost: typical runs taking over a week only achieve a limited $1\%$ accuracy on observables for quenches, and relatively short reachable time scales [4, 5]. Progress is often made by losing the information over the spatial dynamics of the environment [6, 7, 8, 9, 10, 11, 12, 13].
@@
In this Letter, we reveal a crucial feature of quantum impurities: their quantum dynamics involves only a small number of entangled environmental degrees of freedom, called active orbitals, that continuously adapt during the time evolution. The rest of the macroscopic environment remains at each time in a trivial product state (Slater determinant for fermions), making the simulation of systems with thousands of sites now reachable. Entanglement growth is prevented by collecting at each time step the orbital rotation into an evolving unitary frame that is never applied to the state (this would make the problem many-body again), but only to observables. Our main advance is to propose a numerically exact algorithm (with controllable errors below $10^{-6}$) that takes advantage of this few-body structure, leading to a drastic reduction of the computing time (down to only minutes for full relaxation of systems with hundreds of sites), while giving access to real space information on the environment. This algorithm allows us to study the buildup of the dynamical Kondo cloud after a quench, uncovering new aspects of decoherence in a fermionic environment.
@@
Figure 1: MPS bond dimension $\chi_{a}$ as a function of orbital index $a$ for the IRLM with $L=300$ sites at $U/D=-0.5$ and $V/D=0.1$. For ground states (crosses), natural orbitals show a drastic compression w.r.t. real-space and k-space orbitals, because all natural orbitals with $\chi_{a}=1$ are contained in a Slater state. For the dynamics following a quench (circles), an optimal time-dependent orbital rotation remains few-body at the long time $D.t=100$, while $\chi_{a}$ would exceed here $10^{4}$ for real-space and k-space, making the computation unfeasible.
@@
$\hat{H}$ $=$ $U\Big{(}\hat{n}_{0}-\frac{1}{2}\Big{)}\Big{(}\hat{n}_{1}-\frac{1}{2}\Big{)}+\epsilon_{0}(t)\hat{n}_{0}+V(\hat{c}^{\dagger}_{0}\hat{c}^{\phantom{\dagger}}_{1}+\hat{c}^{\dagger}_{1}\hat{c}^{\phantom{\dagger}}_{0})$ (1)
$+\frac{D}{2}\sum_{i=1}^{L-2}(\hat{c}_{i+1}^{\dagger}\hat{c}^{\phantom{\dagger}}_{i}+\mathrm{h.c.}),$
with interaction $U$, hybridization $V$, and hoppings $D/2$. The time-dependent impurity onsite energy will be quenched from $\epsilon_{0}(t)=-\infty$ (full polarisation) at $t<0$, to $\epsilon_{0}(t)=0$ (half-filling) at $t>0$. The IRLM is a very good testbed for quantum impurities since it exhibits Kondo correlations at $U<0$ in the charge sector.
@@
We propose that the physical state at a given time $t$ is given by $\exp[-i\hat{F}(t)]|\psi(t)\rangle$, where $\hat{F}(t)=\sum_{a,b=0}^{L-1}F_{ab}(t)\hat{c}_{a}^{\dagger}\hat{c}^{\phantom{\dagger}}_{b}$ is a quadratic form in terms of fermion operators, parametrized by an hermitian one-body matrix $F_{ab}(t)$ that we call frame. Previous studies have shown disentangling properties of orbital rotations, but were limited by a fixed time-independent frame [23]. Other similar approaches did not take advantage of the further simplifications brought by impurity models [24, Legeza_Rotation]. Picking an optimal frame $F_{ab}(t)$ will keep a very low entanglement in the many-body wavefunction $|\psi(t)\rangle$ for the whole time evolution. This is possible because for impurity models, only a small number of active orbitals $N_{\mathrm{active}}\ll L$ is involved in $|\psi(t)\rangle$, as hinted in Fig. 1. From Thouless theorem [25], the change of frame transforms each orbital to a linear combination as
@@
The algorithm.
To summarize, our numerically exact algorithm operates by simple iterative updates of the frame matrix $F_{ab}(t)$, and by a few-body evolution of the state $|\psi(t)\rangle$ for each time step $dt$, as follows:
@@
Figure 2: Time evolution of the maximum bond dimension $\chi_{\mathrm{max}}$, for the quench dynamics of IRLM at $U/D=0.2$, $V/D=0.1$, $L=200$ sites, and Trotter step $dt=0.1$. Our algorithm using optimized orbital rotation shows bounded entanglement at long times, while the numerical costs diverge for simulations with real space orbitals. Inset: the total number of active orbitals (with $\chi_{a}>1$) is also seen to saturate over time, and dynamics remains few-body.
@@ Results.
We first begin by illustrating the remarkable disentangling feature of our algorithm for a quench of the IRLM. Figure 2 shows the time evolution of the maximum bond dimension $\chi_{\mathrm{max}}=\mathrm{Max}[\chi_{a}]$, comparing calculations using the time-dependent orbital basis or the real space basis. While $\chi_{\mathrm{max}}$ grows unbounded in the latter, it saturates to small values in the former. In addition, the number of active orbitals $n_{\mathrm{active}}$ in the new algorithm remains bounded at long times, making the scheme controlled and efficient.
@@
Figure 3: Impurity charge relaxation toward the equilibrium value $n_{0}(t\to\infty)=1/2$ for the IRLM at $U/D=0.2$, $V/D=0.1$, $L=200$ sites, and Trotter step $dt=0.01$. The dynamics for $U>0$ is coherent, with a leading exponential envelope, and a certain number of zeroes indicated as $t_{1},t_{2},t_{3}$ here. Inset: the first zero $t_{1}$ diverges as $1/U$ for $U\to 0$, while the second zero $t_{2}$ diverges at finite $U_{\mathrm{c}}/D\simeq 0.12$ with a different power law $1/\sqrt{U-U_{\mathrm{c}}}$. The dynamics is thus fully incoherent for $U\leq 0$, and partially coherent for $0<U<U_{c}$.
@@
A crucial test of the accuracy of our new algorithm is given in Fig. 3. Previous time-dependent NRG and DMRG studies [4] for the same quench protocol have shown that errors could not be decreased below $10^{-2}$, preventing a detailed study of charge relaxation at long times. Interestingly, it was predicted [31] that the damping of the charge imbalance at the impurity site, $n_{0}(t)-1/2$, would crossover from a fully coherent regime (with an infinite number of zeroes) at $U>U_{\mathrm{c}}$, to a partially coherent regime (with a finite number of zeroes) for $0<U<U_{\mathrm{c}}$. The fully coherent regime, here displaying three zeroes at times $t_{1}$, $t_{2}$, $t_{3}$, is shown in the main panel of Fig. 3. Extracting these zeroes for many values of $U$ (see inset), we find that the first zero $t_{1}$ survives down to $U\to 0$ and diverges as $1/U$, while the subsequent zero $t_{2}$ diverges at finite $U_{\mathrm{c}}$, confirming for the first time with wavefunction methods the analytical predictions of [31]. In addition, we are able to identify a different power law behavior $t_{2}\propto 1\sqrt{U-U_{\mathrm{c}}}$ governing the disappearance of the second zero. This study surmounts a numerical challenge, because it requires controlling all errors down to $10^{-6}$, as detailed in Sec. III of Ref. [27].
@@
Figure 4: Top: Local charge $\langle\hat{n}_{i}\rangle$ on site $i$ for a long chain of $L=2000$ sites, with $U/D=-0.2$, $V/D=0.1$, and Trotter step $dt=0.1$, showing the ballistic motion of the charge wavepacket at three different times. Middle: charge correlation function $|\langle\hat{n}_{0}\hat{n}_{i}\rangle-\langle\hat{n}_{0}\rangle\langle\hat{n}_{i}\rangle|$ for the same parameters as the top panel, only on even sites. At distances $i\ll Dt$, a charge Kondo cloud binds progressively to the impurity. At distances $i\simeq D.t$, a ballistic peak of correlations between the local charge $\hat{n}_{i}$ in the wavepacket and the impurity charge $\hat{n}_{0}$ is seen, that decays in amplitude as $t\to\infty$, due to loss of entanglement outside the Kondo cloud. Bottom: same plot as the middle panel, but for both even/odd sites, now on a logarithmic scale and only for the time $Dt=1000$. The comparison to the ground state solution (lines) shows proper relaxation to the Kondo cloud for $i\ll Dt$, and also the formation of a dynamical cloud for $i\gtrsim Dt$, despite the fact that the excess charge is vanishing after the wavefront (top panel).
@@
We finally investigate the spatial structures in the environment at long times after the quench for the IRLM at negative $U/D=-0.2$ (Kondo regime, corresponding to a large relaxation time $DT\simeq 100$), for a very large system with $L=2000$ sites. Focusing on the local charge $\langle\hat{n}_{i}\rangle$ (top panel of Fig. 4), we see that the extra density moves away ballistically from the impurity located at site $i=0$, without any damping or dispersion. Spatial Kondo correlations, given by the observable $|\langle\hat{n}_{0}\hat{n}_{i}\rangle-\langle\hat{n}_{0}\rangle\langle\hat{n}_{i}\rangle|$, have been the topic of intense studies for the ground state of impurity models [32]. While Numerical Renormalization Group (NRG) simulations are feasible in and out of equilibrium [33, 34], they require the expensive construction of a double bath for each computed distance in the chain, while our simulation is done with only one bath, and the computation of any observable is single shot. Errors are also maintained below $10^{-4}$ on all sites in our calculation. For the quench, we find several remarkable observations in these spatial correlations. First, at short distances $0<i\ll Dt$ compared to the wavepacket, a Kondo cloud bound to the impurity progressively forms, see middle/bottom panels of Fig. 4 (for linear/logarithmic views). This constitutes a non-trivial check that the relaxation to equilibrium is correctly established in our numerics. Second, we observe a peak of correlations on sites that take part in the wavefront, namely for $i\lesssim D.t$ (middle panel). The emitted charge wavepacket, although it is spread over hundred of sites, maintains a finite entanglement to the impurity. Relaxation to equilibrium is nevertheless reached at $t\to\infty$, since these correlations decay at long times, while the excess charge remains constant in amplitude. Third, we reveal new details on the Kondo-like correlations that emerge in front of the wavepacket for sites at distances $i\gtrsim D.t$ (bottom panel). While “supra-luminous” correlations are expected [34], they remarkably reach very long distances beyond the wavefront, in a region of space where charge neutrality is exactly balanced, and quantitatively differ from the Kondo cloud correlations that are bound to the impurity. A full movie of the dynamics is available here [35].
@@
This work revealed that the dynamics of quantum impurities can be made almost trivial (namely few-body and low entangled) using an optimal time-dependent orbital rotation. This finding gave rise to a numerically exact algorithm that is very simple to implement, faster and more accurate that previous state of the art, and able to cope with macroscopically large systems and arbitrary long times. This advance opens obviously many promising perspectives for the dynamical simulations of interacting quantum systems. For spinful Anderson quantum impurities, the spatio-temporal dynamics of spin-charge separation during protocols such as quenches, or under transport conditions, could be explored with a direct extension of our algorithm. Interaction effects in electronic waveguides [36], often neglected [37], could now be investigated. In the case of superconducting quantum dots, orbital rotations should make the approach of Ref. [38] more systematic, opening the possibility to study the dynamics of interacting Andreev bound states. Since orbital rotation allows us to work with very large baths, it is tempting to apply this approach to finite temperature thermofield simulations [39]. Another promising route concerns real-time quantum impurity solvers for the DMFT equations [2]. Since our algorithm can cope with macroscopically large baths without suffering from the sign problem, accurate multi-orbital calculations in real-frequencies are finally at sight. In addition, extensions to out-of-equilibrium DMFT [40], a topic with fascinating physical questions [41], could pave the way to non-perturbative treatments, still an open problem. Finally, relevant for the effort in simulating superconducting qubits under strong drive [42], the extension of time-dependent orbital rotation to bosons and two-level systems could offer tremendous improvement for the simulations of quantum processing platforms [43].
Acknowledgements. YNF and SF were supported by the PEPR integrated project EPiQ ANR-22-PETQ-0007 part of Plan France 2030. MD was supported by the Deutsche Forschungsgemeinschaft through the cluster of excellence ML4Q (EXC 2004, project-id 390534769) and by the Deutsche Forschungsgemeinschaft through CRC 1639 NuMeriQS (project-id 511713970). We also thank I. Snyman and T. Costi for stimulating discussions.
Supplemental material for “Resolving space-time structures of quantum impurities with time-dependent orbitals”

### v3

# Resolving space-time structures of quantum impurities with a numerically-exact algorithm using few-body revealing

@@ Abstract
We introduce a numerically exact real-time evolution scheme for quantum impurities in a macroscopically large bath. The algorithm is few-body revealing, namely it identifies the electronic orbitals that can be made inactive (in a trivial product state) by a time-dependent orbital rotation. Following a quench, we show that both the number of active orbitals and their associated matrix product state bond dimensions saturate to small values, leading to an algorithm dramatically more accurate and faster than the state of the art. We are thus able to follow the dynamics for thousands of fermions, up to the long-time stationary regime, and to study subtle aspects of quantum relaxation in the spatio-temporal domain, such as the emergence of entanglement structures in the Kondo screening cloud.
@@
The many-body dynamics of generic quantum systems currently challenges numerical simulations, as exemplified by the unbounded entanglement growth in time of matrix product states (MPS). This situation also plagues quantum impurity problems, “simpler” systems still comprised of a macroscopic number of particles (bosons or fermions), where the interaction is restricted to a subset of degrees of freedom [1]. The interest in impurity models, originating in Kondo physics, has not waned over the years, since they are the cornerstone of the Dynamical Mean-Field Theory (DMFT) for correlated electrons [2]. Open quantum systems, in which qubits exchange energy and information with an environment, constitute also a very important class of impurity models [3]. However, in most cases, simulating the real-time non-perturbative dynamics of impurity models comes with tremendous numerical cost: typical runs taking over a week only achieve a limited $1\%$ accuracy on observables for quenches, and relatively short reachable time scales [4, 5]. Progress is often made by losing the information over the spatial dynamics of the environment [6, 7, 8, 9, 10, 11, 12, 13, 14].
@@
We pose in this Letter the idea of few-body revealing for quantum impurities models, based on a new finding that their quench dynamics involves only a small number of entangled environmental degrees of freedom, called active orbitals, that continuously adapt during the time evolution. The rest of the macroscopic environment remains at each time in a trivial, yet evolving, product state (Slater determinant for fermions), making the simulation of systems with thousands of sites now reachable. Entanglement growth is prevented by collecting at each time step the orbital rotation into an evolving unitary frame that is never applied to the state (this would make the problem many-body again), but only to observables. Our main advance is to propose a numerically exact algorithm (with controllable errors below $10^{-6}$) that takes advantage of this few-body structure, leading to a drastic reduction of the computing time (down to only minutes for full relaxation of systems with hundreds of sites), while giving access to real space information on the environment. This algorithm allows us to study the buildup of the dynamical Kondo cloud after a quench, uncovering new aspects of decoherence in a fermionic environment.
@@
Figure 1: Demonstrating few-body revealing from the MPS bond dimension $\chi_{a}$ as a function of orbital index $a$, for the IRLM with $L=300$ sites at $U/D=-0.5$ and $V/D=0.1$. For ground states (crosses), natural orbitals show a drastic compression w.r.t. real-space and k-space orbitals, because all natural orbitals with $\chi_{a}=1$ are contained in a Slater state. For the dynamics following a quench (circles), an optimal orbital rotation leaves the state vector $|\psi(t)\rangle$ few-body at the long time $Dt=100$, while $\chi_{a}$ would exceed here $10^{4}$ for real-space and k-space, making the computation unfeasible.
@@
$\hat{H}$ $=$ $U\Big{(}\hat{n}_{0}-\frac{1}{2}\Big{)}\Big{(}\hat{n}_{1}-\frac{1}{2}\Big{)}+\epsilon_{0}(t)(\hat{n}_{0}-\hat{n}_{1})$ (1)
$+V(\hat{c}^{\dagger}_{0}\hat{c}^{\phantom{\dagger}}_{1}+\hat{c}^{\dagger}_{1}\hat{c}^{\phantom{\dagger}}_{0})+\frac{D}{2}\sum_{i=1}^{L-2}(\hat{c}_{i+1}^{\dagger}\hat{c}^{\phantom{\dagger}}_{i}+\mathrm{h.c.}),$
with interaction $U$, hybridization $V$, and hoppings $D/2$. The time-dependent impurity onsite energy will be quenched from $\epsilon_{0}(t)=-\infty$ at $t<0$ (so that sites 00 and $1$ are initially decoupled from the bath), to $\epsilon_{0}(t)=0$ at $t>0$. The IRLM is a very good testbed for quantum impurities since it exhibits Kondo correlations at $U<0$ in the charge sector.
@@
We propose that the physical state at a given time $t$ is given by $\exp[-i\hat{F}(t)]|\psi(t)\rangle$, where $\hat{F}(t)=\sum_{a,b=0}^{L-1}F_{ab}(t)\hat{c}_{a}^{\dagger}\hat{c}^{\phantom{\dagger}}_{b}$ is a quadratic form in terms of fermion operators, parametrized by an hermitian one-body matrix $F_{ab}(t)$ that we call frame. Previous studies have shown disentangling properties of orbital rotations [24, 25, 26, 27, 28, 29, 30], but were limited by a fixed time-independent frame, or did not take full advantage of the simplifications brought by impurity models. Picking an optimal frame $F_{ab}(t)$ will keep a very low entanglement in the many-body wavefunction $|\psi(t)\rangle$ for the whole time evolution. This is possible because for the quench dynamics of impurity models, only a small number of active orbitals $N_{\mathrm{active}}\ll L$ is involved in $|\psi(t)\rangle$, as hinted in Fig. 1. From Thouless theorem [31], the change of frame transforms each orbital to a linear combination as
@@
The few-body revealing algorithm.
To summarize, our algorithm operates by simple iterative updates of the frame matrix $F_{ab}(t)$, and by a few-body evolution of the state $|\psi(t)\rangle$ for each time step $dt$, as follows:
@@
Figure 2: Time-evolution of few-body revealing, seen from the maximum bond dimension $\chi_{\mathrm{max}}$, for the quench dynamics of IRLM at $U/D=0.2$, $V/D=0.1$, $L=200$ sites, and Trotter step $dt=0.1$. Our algorithm using time-dependent orbital rotation shows bounded entanglement at long times, while the numerical costs diverge for simulations with real space orbitals. Inset: the total number of active orbitals (with $\chi_{a}>1$) is also seen to saturate over time, and dynamics remains few-body.
@@ Results.
We first begin by illustrating the remarkable disentangling feature of our few-body revealing algorithm for a quench of the IRLM. Figure 2 shows the time evolution of the maximum bond dimension $\chi_{\mathrm{max}}=\mathrm{Max}[\chi_{a}]$, comparing calculations using the time-dependent orbital basis or the real space basis. While $\chi_{\mathrm{max}}$ grows unbounded in the latter, it saturates to small values in the former. In addition, the number of active orbitals $N_{\mathrm{active}}$ in the new algorithm remains bounded at long times, making the scheme controlled and efficient.
@@
Figure 3: Impurity charge relaxation toward the equilibrium value $n_{0}(t\to\infty)=1/2$ for the IRLM at $U/D=0.2$, $V/D=0.1$, $L=200$ sites, and Trotter step $dt=0.01$. The dynamics for $U>0$ is coherent, with a leading exponential envelope, and a certain number of zeros indicated as $t_{1},t_{2},t_{3}$ here. Inset: the first zero $t_{1}$ diverges as $1/U$ for $U\to 0$ (in agreement with analytical results extracted from Ref. [37]), while the second zero $t_{2}$ is found to diverge at finite $U_{\mathrm{c}}^{t_{2}}/D\simeq 0.12$ with a different power law $(U-U_{\mathrm{c}}^{t_{2}})^{-\frac{1}{2}}$.
@@
A crucial test of the accuracy of our new algorithm is given in Fig. 3. Previous time-dependent NRG and DMRG studies [4] for the same quench protocol have shown that errors could not be decreased below $10^{-2}$, preventing a detailed study of charge relaxation at long times. Interestingly, it was predicted [37] that the damping of the charge imbalance at the impurity site, $n_{0}(t)-1/2$, would crossover from a fully coherent regime (with an infinite number of zeros) at $U>U_{\mathrm{c}}$, to a partially coherent regime (with a finite number of zeros) for $0<U<U_{\mathrm{c}}$. Such coherent regime, here displaying three zeros at times $t_{1}$, $t_{2}$, $t_{3}$, is shown in the main panel of Fig. 3. We extracted these zeros for many $U$ values (see inset), and we observe that the first zero $t_{1}$ disappears at $U\to 0$ as expected from the existence of the incoherent regime at $U\leq 0$ [38], and diverges as $1/U$, in agreement with the analytics of Ref. [37]. Interestingly, we find that the subsequent zero $t_{2}$ diverges at a finite $U_{\mathrm{c}}^{t_{2}}$, a new observation implying that relaxation occurs via a unique zero $t_{1}$ for $0<U<U_{\mathrm{c}}^{t_{2}}$. This is compatible with (yet clearly beyond) the predictions of Ref. [37]. In addition, we are able to identify a different power law behavior $t_{2}\propto(U-U_{\mathrm{c}}^{t_{2}})^{-\frac{1}{2}}$ governing the disappearance of the second zero. Our study surmounts a numerical challenge, because it requires controlling all errors down to $10^{-6}$, as detailed in Sec. III of Ref. [33].
@@
Figure 4: Top: Local charge $\langle\hat{n}_{i}\rangle$ on site $i$ for a long chain of $L=2000$ sites, with $U/D=-0.2$, $V/D=0.1$, and Trotter step $dt=0.1$, showing the ballistic motion of the charge wavepacket at three different times. Middle: charge correlation function $|\langle\hat{n}_{0}\hat{n}_{i}\rangle-\langle\hat{n}_{0}\rangle\langle\hat{n}_{i}\rangle|$ for the same parameters as the top panel, only on even sites. At distances $i\ll Dt$, a charge Kondo cloud binds progressively to the impurity. At distances $i\simeq Dt$, a ballistic peak of correlations between the local charge $\hat{n}_{i}$ in the wavepacket and the impurity charge $\hat{n}_{0}$ is seen, that decays in amplitude as $t\to\infty$, due to loss of entanglement outside the Kondo cloud. Bottom: same plot as the middle panel, but for both even/odd sites, now on a logarithmic scale and only for the time $Dt=1000$. The comparison to the ground state solution (lines) shows proper relaxation to the Kondo cloud for $i\ll Dt$, and also the formation of a dynamical cloud for $i\gtrsim Dt$, despite the fact that the excess charge is vanishing after the wavefront (top panel).
@@
We finally investigate the spatial structures in the environment at long times after the quench for the IRLM at negative $U/D=-0.2$ (Kondo regime, corresponding to a large relaxation time $Dt\simeq 100$), for a very large system with $L=2000$ sites. Such dynamics can be easily computed in the non-interacting case [39, 33], but is very challenging in presence of interactions. Focusing on the local charge $\langle\hat{n}_{i}\rangle$ (top panel of Fig. 4), we see that the extra density moves away ballistically from the impurity located at site $i=0$, without any damping or dispersion. Spatial Kondo correlations, given by the observable $|\langle\hat{n}_{0}\hat{n}_{i}\rangle-\langle\hat{n}_{0}\rangle\langle\hat{n}_{i}\rangle|$, have been the topic of intense studies for the ground state of impurity models [40]. While Numerical Renormalization Group (NRG) simulations are feasible in and out of equilibrium [41, 42], they require the expensive construction of a double bath for each computed distance in the chain, while our simulation is done with only one bath, and the computation of any observable is single shot. Errors are also maintained below $10^{-4}$ on all sites in our calculation. For the quench, we find several remarkable observations in these spatial correlations. First, at short distances $0<i\ll Dt$ compared to the wavepacket, a Kondo cloud bound to the impurity progressively forms, see middle/bottom panels of Fig. 4 (for linear/logarithmic views). This constitutes a non-trivial check that the relaxation to equilibrium is correctly established in our numerics. Second, we observe a peak of correlations on sites that take part in the wavefront, namely for $i\lesssim Dt$ (middle panel). The emitted charge wavepacket, although it is spread over hundred of sites, maintains a finite entanglement to the impurity. Relaxation to equilibrium is nevertheless reached at $t\to\infty$, since these correlations decay at long times, while the excess charge remains constant in amplitude. Third, we reveal new details on the Kondo-like correlations that emerge in front of the wavepacket for sites at distances $i\gtrsim$ (bottom panel). While “supra-luminous” correlations are expected [42], they remarkably reach very long distances beyond the wavefront, in a region of space where charge neutrality is exactly balanced, and quantitatively differ from the Kondo cloud correlations that are bound to the impurity. A full movie of the dynamics is available here [43].
@@
Using optimal time-dependent orbital rotations, we showed that few-body revealing makes the quench dynamics of quantum impurities almost trivial (namely few-body and low entangled). This finding gives rise to a numerically exact algorithm that is very simple to implement, faster and more accurate that previous state of the art, and able to cope with macroscopically large systems and long times. This advance opens obviously many promising perspectives for the dynamical simulations of interacting quantum systems. For spinful Anderson quantum impurities, the spatio-temporal dynamics of spin-charge separation could be explored, possibly at finite temperature with a thermofield approach [28]. Generalizing orbital rotations to superconducting quantum dots [44] should also open the possibility to study the dynamics of interacting Andreev bound states. Interaction effects in electronic waveguides [45], often neglected [46], could also be investigated. Another promising route concerns real-time quantum impurity solvers [25] for the DMFT equations [2], especially in the multi-orbital case [47], and in the context of out-of-equilibrium physics [48], a fascinating and challenging topic [49]. Finally, extensions of time-dependent orbital rotation to bosons and two-level systems could help the simulation of driven superconducting qubits [50] and quantum processing platforms [51].
Acknowledgements. YNF and SF were supported by the PEPR integrated project EPiQ ANR-22-PETQ-0007 part of Plan France 2030. MD was supported by the Deutsche Forschungsgemeinschaft through the cluster of excellence ML4Q (EXC 2004, project-id 390534769) and by the Deutsche Forschungsgemeinschaft through CRC 1639 NuMeriQS (project-id 511713970). We also thank T.Costi, V. Meden, and I. Snyman for stimulating discussions.
Supplemental material for “Resolving space-time structures of quantum impurities with a numerically-exact algorithm using few-body revealing”